# Exact predicates, exact constructions and combinatorics

### v1

@@ 2.1. Mesh co-refinement
Mesh co-refinement takes as input a set of triangulated surface. No assumption is made regarding the structure of this input. It can be a triangle soup made of disconnected triangles. It can have co-planar facets. Duplicated vertices are pre-detected and merged using spatial sorting. Duplicated facets are then pre-detected and discarded, using lexicographic sort. Then, the different substeps of the algorithm, outlined in Figure 3, are as follows:
@@ 2.1.2. Computing the intersection between two triangles
To tackle this problem, we will write a function that computes a combinatorial representation of the intersection. Each triangle $t$ (resp. $t^{\prime}$) can be seen as a simplicial set $\Sigma_{t}$ with 7 simplices: each triangle has three vertices $V_{1},V_{2},V_{3}$, three edges $E_{1}=(P_{2},P_{3})$, $E_{2}=(P_{3},P_{1})$, $E_{3}=(P_{1},P_{2})$, and the whole triangle $T=(P_{1},P_{2},P_{3})$. We consider that each simplex is embedded as an open set. In other words, each edge is embedded as a segment minus the extremities, and $T$ is embedded as the triangle minus its border, hence each point of the triangle is contained by exactly one simplex. For each intersection point (in red in Figure 4) we will output the unique pair of simplices $\sigma,\sigma^{\prime}$ whose embeddings contain the point, where $\sigma\in\Sigma_{t}=\{P_{1},P_{2},P_{3},E_{1},E_{2},E_{3},T\}$ and $\sigma^{\prime}\in\Sigma_{t^{\prime}}=\{P_{1}^{\prime},P_{2}^{\prime},P_{3}^{\prime},E_{1}^{\prime},E_{2}^{\prime},E_{3}^{\prime},T^{\prime}\}$. In practice, in an implementation, the simplices can be encoded as integers, or enums. From this description, one can imagine the following (generic, naive) triangle-triangle intersection algorithm, that tests all the $7\times 7$ possible intersections between the simplices of $t$ and $t^{\prime}$:
@@
$\begin{array}[]{ll}&\mbox{edge\_triangle()}\\ &\mbox{\bf input:}\quad\mbox{an edge }E\mbox{ and a triangle }T\\ &\mbox{\bf output:}\quad\mbox{a list }I\mbox{ of simplices pairs }(\sigma,\sigma^{\prime})\\ &\quad\quad\mbox{ that define the intersection points}\\ \hline\cr\\ (1)&(P_{1},P_{2},P_{3})\leftarrow\mbox{vertices}(T)\\ (2)&(Q_{1},Q_{2})\leftarrow\mbox{vertices}(E)\\ (3)&s_{1}\leftarrow\mbox{orient3d}(P_{1},P_{2},P_{3},Q_{1})\\ (4)&s_{2}\leftarrow\mbox{orient3d}(P_{1},P_{2},P_{3},Q_{2})\\ (5)&\mbox{\bf if }s_{1}\times s_{2}>0\ \mbox{\bf then }\mbox{\bf return }\emptyset\\ (6)&\mbox{\bf if }s_{1}=0\ \mbox{\bf and }s_{2}=0\ \mbox{\bf then }\mbox{\bf return}\mbox{edge\_triangle\_2D}(E,T)\\ (7)&o_{1}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{1},P_{2})\\ (8)&o_{2}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{2},P_{3})\\ (9)&o_{3}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{3},P_{1})\\ (10)&\mbox{\bf if }o_{1}\times o_{2}<0\ \mbox{\bf or }o_{2}\times o_{3}<0\ \mbox{\bf or }o_{3}\times o_{1}<0\ \mbox{\bf then }\mbox{\bf return }\emptyset\\ (11)&R_{1}\leftarrow\mbox{region}(T,o_{1},o_{2},o_{3})\\ (12)&R_{2}\leftarrow\mbox{region}(E,s_{1},s_{2})\\ (13)&\mbox{\bf return }(R_{2},R_{1})\end{array}$
@@
The number of invocations to the orient_3d predicate remains higher than the optimal, because our algorithm naively computes intersection between sub-simplices without taking into account more global information. To avoid doing the same computations several times, we use a “predicate cache”, indexed by the sorted list of its four arguments, and changing the sign in function of the parity of the actual order of the four arguments (orient_3d is a determinant, hence permutting its columns changes the sign depending on the parity of the permutation).
@@ 2.1.3. Constrained Delaunay Triangulation
In addition to the triangulation beeing constrained, it is interesting to require it to be a constrained Delaunay triangulation for two reasons:
  • With Delaunay, the quality of the mesh is “not too bad”, because it maximizes the smallest angle. We say here “not too bad” because in general, even with Delaunay, intersection meshes contain small angles and cannot be directly used in numerical simulation without some re-meshing / post-processing. However, it is always good to have a starting point that does not have too many triangles with very small angles;
  • the Delaunay triangulation is unique, which is an interesting property when the intersecting meshes have overlapping coplanar facets. This property ensures that the same zone will be meshed with the same triangles (one only needs to filter-out the duplicated triangles). There exists other method to solve the problem with non-unique triangulations, based on a cavity-remeshing operator with linear complexity (Livesu et al., 2022), but this require maintaining a list of polygons together with the triangulations, and identifying the duplicated polygons, using a more complicated data structure. In other words, this means pushing the difficulties into the combinatorial data structure. We prefer to keep them in the predicates, because predicates are concentrated in a small portion of the code, easier to maintain and to debug.
@@
There are several implementations of a constrained Delaunay triangulation available, such as Shewchuk’s Triangle (Shewchuk, 1996b) and CGAL (The CGAL Project, 2023a). We chose not to use them for several reasons. First, Triangle has hardwired predicates (whereas we need to plug special ones, adapted to points that come from intersections), and it has global variables, preventing it to be used in a multithreaded context. CGAL can be completely parameterized through a template mechanism. However, by default, it internally uses pointer-based data structures that do dynamic allocations, which has an impact of performance in a multithreaded context.
@@
$\begin{array}[]{ll}&\mbox{constrained\_Delaunay\_triangulation}\\ &\mbox{\bf input:}\\ &\quad\mbox{a triangle }t_{0}=(p_{0},p_{1},p_{2})\\ &\quad\mbox{a list of vertices }p_{i},i=3\ldots N_{v}\mbox{ inside }t\\ &\quad\mbox{a list of edges }E_{k}=(i_{k},j_{k}),k=1\ldots N_{e}\\ &\mbox{\bf output:}\\ &\quad\mbox{the Delaunay triangulation of the }p_{i}\mbox{'s}\\ &\quad\quad\mbox{constrained by the }E_{k}\mbox{'s and by }t\mbox{'s edges}.\\ \hline\cr\\ (1)&\mbox{\bf for }i=1\ \mbox{\bf to }N_{v}\\ (2)&\quad\mbox{find the triangle }t\mbox{ that contains }p_{i}\\ (3)&\quad\mbox{insert }p_{i}\mbox{ into }t\\ (4)&\quad\mbox{push the three triangle edges opposite to }p_{i}\mbox{ onto }S\\ (5)&\quad\mbox{Delaunayize\_vertex\_neighbors}(p_{i},S)\\ (6)&\mbox{\bf end }\\ (7)&\mbox{\bf for }k=1\ \mbox{\bf to }N_{e}\\ (8)&\quad\mbox{enqueue the edges intersecting }(i_{k},j_{k})\mbox{ onto }Q\\ (9)&\quad N\leftarrow\mbox{constrain\_edges}(i_{k},j_{k},Q)\\ (10)&\quad\mbox{Delaunayize\_new\_edges}(N)\\ (11)&\mbox{\bf end }\end{array}$
@@
(lines (1) to (6)) inserts the vertices one by one in the triangulation, by first locating the triangle $t$ that contains the point $p_{i}$. Then it splits this triangle into three (note that $p_{i}$ can be exactly located on an edge, then the two triangles that share that edge are split into two, for a total of four new triangles). Then the Delaunay condition is restored by the Delaunayize_vertex_neighbors() function. This function recursively flips the edges that violate the Delaunay condition and pushes the new triangle on the stack $S$ until the stack is empty. The reader is referred to the original article (Sloan, 1992) for more details.
@@
(lines (7) to (11)) inserts the constraints one by one in the triangulation. The first step (line (8)) detects the edges that have an intersection with the constraint. This is done by “walking the triangulation” along the edge, one triangle at a time, and testing for each triangle two edges (the third one is the one we came from). If one of the intersected edge is a constraint, then this means we have detected a triple point, like in Figure 1-C, where the intersection of edges $e_{1}$ and $e_{2}$ generate a new vertex $I$. The intersected edges are pushed to a queue $Q$. Then the function constrain_edges() processes each “flippable” edge of the queue until the queue is empty. By “flippable”, we mean that the two triangles adjacent to the edge form a convex polygon. It can be proven that this process converges. Each time an edge is flipped, the corresponding triangles are saved in a list $N$ of “new” triangles, finally processed by the Delaunayize_new_edges() function that flips edges until the Delaunay condition is satisfied everywhere.
@@
    first, we are going to compute a huge number of constrained Delaunay triangulations, in each individual triangle that has intersections. To keep performance acceptable, we are going to construct them in parallel. Dynamic memory allocation is a serious obstacle to efficient parallel code, because there is a global lock associated with the malloc() function (or the new() operator), so our data structure will be solely composed of std::vector’s allocated once for all666except at the beginning when they will grow as needed, and later, they never shrink.. We also need a data structure for the stack of triangles $S$ and the queue of triangles $Q$. We use a doubly connected list, represented as two additional std::vector’s that store the forward and backward link;
@@
The algorithm to restore de Delaunay condition around new vertices and around new edges is as in the classical implementations of the Bowyer-Watson algorithm. The algorithm to constrain the edges deserves more details, because the way the edges are systematically manipulated through triangles reveals an optimization that avoids most invocations to the ${\tt orient\_2d}$ predicate. The classical version is as follows:
@@
The segment-segment intersection test in line (5) that normally requires up to 4 invocations of the ${\tt orient\_2d}$ predicate can be replaced with a single invocation and some combinatorics. Remember that each edge $(v_{1},v_{2})$ is systematically manipulated through a triangle $(v_{0},v_{1},v_{2})$, rotated in such a way that its vertex $v_{0}$ is opposite to the edge under consideration. Then, as shown in Figure 7, there are four different configurations for the pair of triangles $t_{1}=(v_{0},v_{1},v_{2})$ and $t_{2}$ adjacent to $t_{1}$ along $(v_{1},v_{2})$, depending on whether $t_{2}$ is in the queue $Q$ of intersected edges777that is tested in O(1) using a per-triangle array of marks., depending on ${\tt orient\_2d(i,j,v_{0})}$ and depending on whether $t_{2}$’s first vertex is $v_{1}$ or $v_{2}$:
@@
So now the question is how to implement these three functions given that some of our vertices are given as intersections between segments and triangles. The strategy here is to keep the algorithm as near as possible to “simply implement the math”. In other words, it means we are pushing most of the difficulties towards the geometric kernel (predicates and constructions) to keep the overall structure of the algorithm simple (nearly a verbatim copy of the textbook algorithm). So we are going to compute the coordinates of the intersection points explicitely. Since these coordinates are rational fractions, and since computations for $x,y$ and $z$ are inter-related, it is reasonable to represent the intersection points in homogeneous coordinates $\hat{{\bf p}}=[x\ y\ z\ w]$ that corresponds to a 3D point ${\bf p}=[x/w\ y/w\ z/w]$. Each individual $x,y,z,w$ coordinate is represented in an exact number type that exactly implements addition, subtraction and multiplication. We shall see later (§2.4 p.2.4) two alternatives to implement exact number types, and how to implement all the predicates that we need.
The input of the constrained Delaunay triangulation is a triangle $t_{1}$ and the list of segments to be inserted into $t_{1}$. Each segment’s extremity is encoded symbolically, as a triple $\sigma_{1},t_{2},\sigma_{2}$, indicating the location of the intersection within $t_{1}$, the other facet $t_{2}$ and the location of the intersection within $t_{2}$. Note that there can be also individual points (for instance, when a vertex is exactly located in a facet). The first task to do is computing the coordinates of each intersection. Depending on the nature of $\sigma_{1}$ and $\sigma_{2}$, there are three different cases, plus an additional case for intersecting constraints:
@@
Let ${\bf q}_{1}$ and ${\bf q}_{2}$ denote the extremities of the edge and let ${\bf p}_{1}$,${\bf p}_{2}$ and ${\bf p}_{3}$ denote the vertices of the triangle. We are in 3D if ${\bf q}_{1}$ and ${\bf q}_{2}$ are not both in the supporting plane of ${\bf p}_{1}$,${\bf p}_{2},{\bf p}_{3}$ (in other words, at least one of orient_3d$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf q}_{1})$ and orient_3d$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf q}_{2})$ is non-zero). Note that if we are in 3D, the edge $e_{1}$cap edge $e_{2}$ configuration can be seen as a particular case of the edge $e_{1}$cap triangle $t_{2}$ case, by replacing $e_{2}$ by the triangle $t_{2}$ it comes from. At the previous step, we already determined that the intersection exists, so to compute its coordinate we just need to compute a line-plane intersection, without needing to check that the intersection is in the segment and in the triangle. The intersection point ${\bf I}$ belongs to the line:

### v2

@@ 2.1. Mesh co-refinement
Mesh co-refinement takes as input a set of triangulated surfaces. No assumption is made regarding the structure of this input. It can be a triangle soup made of disconnected triangles. It can have co-planar facets. Duplicated vertices are pre-detected and merged using lexicographic sorting. Duplicated facets are then pre-detected and discarded, using lexicographic sort. Then, the different substeps of the algorithm, outlined in Figure 3, are as follows:
@@ 2.1.2. Computing the intersection between two triangles
To tackle this problem, we will write a function that computes a combinatorial representation of the intersection. Each triangle $t$ (resp. $t^{\prime}$) can be seen as a simplicial set $\Sigma_{t}$ with 7 simplices: each triangle has three vertices $V_{1},V_{2},V_{3}$, three edges $E_{1}=(P_{2},P_{3})$, $E_{2}=(P_{3},P_{1})$, $E_{3}=(P_{1},P_{2})$, and the whole triangle $T=(P_{1},P_{2},P_{3})$.
We consider that each simplex is embedded as an open set. In other words, each edge is embedded as a segment minus the extremities, and $T$ is embedded as the triangle minus its border, hence each point of the triangle is contained by exactly one simplex. For each intersection point (in red in Figure 4) we will output the unique pair of simplices $\sigma,\sigma^{\prime}$ whose embeddings contain the point, where $\sigma\in\Sigma_{t}=\{P_{1},P_{2},P_{3},E_{1},E_{2},E_{3},T\}$ and $\sigma^{\prime}\in\Sigma_{t^{\prime}}=\{P_{1}^{\prime},P_{2}^{\prime},P_{3}^{\prime},E_{1}^{\prime},E_{2}^{\prime},E_{3}^{\prime},T^{\prime}\}$. In practice, in an implementation, the simplices can be encoded as integers, or enums. From this description, one can imagine the following (generic, naive) triangle-triangle intersection algorithm, that tests all the $7\times 7$ possible intersections between the simplices of $t$ and $t^{\prime}$:
@@
$\begin{array}[]{ll}&\mbox{edge\_triangle()}\\ &\mbox{\bf input:}\quad\mbox{an edge }E\mbox{ and a triangle }T\\ &\quad\mbox{{\color[rgb]{0,0,0}(both considered as closed sets, with extremities and edges)}}\\ &\mbox{\bf output:}\quad\mbox{a list }I\mbox{ of simplices pairs }(\sigma,\sigma^{\prime})\\ &\quad\quad\mbox{ that define the intersection points}\\ \hline\cr\\ (1)&(P_{1},P_{2},P_{3})\leftarrow\mbox{vertices}(T)\\ (2)&(Q_{1},Q_{2})\leftarrow\mbox{vertices}(E)\\ (3)&s_{1}\leftarrow\mbox{orient3d}(P_{1},P_{2},P_{3},Q_{1})\\ (4)&s_{2}\leftarrow\mbox{orient3d}(P_{1},P_{2},P_{3},Q_{2})\\ (5)&\mbox{\bf if }s_{1}\times s_{2}>0\ \mbox{\bf then }\mbox{\bf return }\emptyset\\ (6)&\mbox{\bf if }s_{1}=0\ \mbox{\bf and }s_{2}=0\ \mbox{\bf then }\mbox{\bf return}\mbox{edge\_triangle\_2D}(E,T)\\ (7)&o_{1}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{1},P_{2})\\ (8)&o_{2}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{2},P_{3})\\ (9)&o_{3}\leftarrow\mbox{orient3d}(Q_{1},Q_{2},P_{3},P_{1})\\ (10)&\mbox{\bf if }o_{1}\times o_{2}<0\ \mbox{\bf or }o_{2}\times o_{3}<0\ \mbox{\bf or }o_{3}\times o_{1}<0\ \mbox{\bf then }\mbox{\bf return }\emptyset\\ (11)&R_{1}\leftarrow\mbox{region}(T,o_{1},o_{2},o_{3})\\ (12)&R_{2}\leftarrow\mbox{region}(E,s_{1},s_{2})\\ (13)&\mbox{\bf return }(R_{2},R_{1})\end{array}$
@@
The number of invocations to the orient_3d predicate remains higher than the optimal, because our algorithm naively computes intersection between sub-simplices without taking into account more global information. To avoid doing the same computations several times, I use a “predicate cache” (a similar technique is used in (Cherchi et al., 2022)). A predicate cache is a table that maps predicate’s input to the resulting sign. The key used to index the table is the list of the sorted indices of the input points. When querying the predicate, if present in the table, the stored sign is flipped depending on the parity of the actual order of the four arguments (orient_3d is a determinant, hence permutting its columns changes the sign depending on the parity of the permutation).
@@ 2.1.3. Constrained Delaunay Triangulation
In addition to the triangulation being constrained, it is interesting to require it to be a constrained Delaunay triangulation for two reasons:
  • With Delaunay, the quality of the mesh is “not too bad”, because it maximizes the smallest angle. I say here “not too bad” because in general, even with Delaunay, intersection meshes contain small angles and cannot be directly used in numerical simulation without some re-meshing / post-processing. However, it is always good to have a starting point that does not have too many triangles with very small angles;
  • the Delaunay triangulation is unique, which is an interesting property when the intersecting meshes have overlapping coplanar facets. This property ensures that the same zone will be meshed with the same triangles (one only needs to filter-out the duplicated triangles). There exists other method to solve the problem with non-unique triangulations, based on a cavity-remeshing operator with linear complexity (Livesu et al., 2022), but this require maintaining a list of polygons together with the triangulations, and identifying the duplicated polygons, using a more complicated data structure. In other words, this means pushing the difficulties into the combinatorial data structure. I prefer to keep them in the predicates, because predicates are concentrated in a small portion of the code, easier to maintain and to debug.
@@
There are several implementations of a constrained Delaunay triangulation available, such as Shewchuk’s Triangle (Shewchuk, 1996b) and CGAL (The CGAL Project, 2023a). I chose not to use them for several reasons. First, Triangle has hardwired predicates (whereas we need to plug special ones, adapted to points that come from intersections), and it has global variables, preventing it to be used in a multithreaded context. CGAL can be completely parameterized through a template mechanism. However, by default, it internally uses pointer-based data structures that do dynamic allocations, which has an impact of performance in a multithreaded context.
@@
$\begin{array}[]{ll}&\mbox{constrained\_Delaunay\_triangulation}\\ &\mbox{\bf input:}\\ &\quad\mbox{a triangle }t_{0}=(p_{0},p_{1},p_{2})\\ &\quad\mbox{a list of vertices }p_{i},i=3\ldots N_{v}\mbox{ inside }t\\ &\quad\mbox{a list of edges }E_{k}=(i_{k},j_{k}),k=1\ldots N_{e}\\ &\mbox{\bf output:}\\ &\quad\mbox{the Delaunay triangulation of the }p_{i}\mbox{'s}\\ &\quad\quad\mbox{constrained by the }E_{k}\mbox{'s and by }t\mbox{'s edges}.\\ \hline\cr\\ (1)&\mbox{\bf for }i=1\ \mbox{\bf to }N_{v}\\ (2)&\quad\mbox{find the triangle }t\mbox{ that contains }p_{i}\\ (3)&\quad\mbox{insert }p_{i}\mbox{ into }t\\ (4)&\quad\mbox{push the three triangle edges opposite to }p_{i}\mbox{ onto }S\\ (5)&\quad\mbox{Delaunize\_vertex\_neighbors}(p_{i},S)\\ (6)&\mbox{\bf end }\\ (7)&\mbox{\bf for }k=1\ \mbox{\bf to }N_{e}\\ (8)&\quad\mbox{enqueue the edges intersecting }(i_{k},j_{k})\mbox{ onto }Q\\ (9)&\quad N\leftarrow\mbox{constrain\_edges}(i_{k},j_{k},Q)\\ (10)&\quad\mbox{Delaunize\_new\_edges}(N)\\ (11)&\mbox{\bf end }\end{array}$
@@
(lines (1) to (6)) inserts the vertices one by one in the triangulation, by first locating the triangle $t$ that contains the point $p_{i}$. Then it splits this triangle into three (note that $p_{i}$ can be exactly located on an edge, then the two triangles that share that edge are split into two, for a total of four new triangles). Then the Delaunay condition is restored by the Delaunize_vertex_neighbors() function. This function recursively flips the edges that violate the Delaunay condition and pushes the new triangle on the stack $S$ until the stack is empty. The reader is referred to the original article (Sloan, 1992) for more details.
@@
(lines (7) to (11)) inserts the constraints one by one in the triangulation. The first step (line (8)) detects the edges that have an intersection with the constraint. This is done by “walking the triangulation” along the edge, one triangle at a time, and testing for each triangle two edges (the third one is the one we came from). If one of the intersected edge is a constraint, then this means we have detected a triple point, like in Figure 1-C, where the intersection of edges $e_{1}$ and $e_{2}$ generate a new vertex $I$. The intersected edges are pushed to a queue $Q$. Then the function constrain_edges() processes each “flippable” edge of the queue until the queue is empty. By “flippable”, we mean that the two triangles adjacent to the edge form a convex polygon. It can be proven that this process converges. Each time an edge is flipped, the corresponding triangles are saved in a list $N$ of “new” triangles, finally processed by the Delaunize_new_edges() function that flips edges until the Delaunay condition is satisfied everywhere.
@@
    first, we are going to compute a huge number of constrained Delaunay triangulations, in each individual triangle that has intersections. To keep performance acceptable, we are going to construct them in parallel. Dynamic memory allocation is a serious obstacle to efficient parallel code, because there is a global lock associated with the malloc() function (or the new() operator), so our data structure will be solely composed of std::vector’s allocated once for all777except at the beginning when they will grow as needed, and later, they never shrink.. We also need a data structure for the stack of triangles $S$ and the queue of triangles $Q$. I use a doubly connected list, represented as two additional std::vector’s that store the forward and backward link;
@@
The algorithm to restore the Delaunay condition around new vertices and around new edges is as in the classical implementations of the Bowyer-Watson algorithm. The algorithm to constrain the edges deserves more details, because the way the edges are systematically manipulated through triangles reveals an optimization that avoids most invocations to the ${\tt orient\_2d}$ predicate. The classical version is as follows:
@@
I make here a simple yet important observation that spares 3/4 of the calls to the orient_2d predicate: The segment-segment intersection test in line (5) that normally requires up to 4 invocations of the ${\tt orient\_2d}$ predicate can be replaced with a single invocation and some combinatorics. Remember that each edge $(v_{1},v_{2})$ is systematically manipulated through a triangle $(v_{0},v_{1},v_{2})$, rotated in such a way that its vertex $v_{0}$ is opposite to the edge under consideration. Then, as shown in Figure 7, there are four different configurations for the pair of triangles $t_{1}=(v_{0},v_{1},v_{2})$ and $t_{2}$ adjacent to $t_{1}$ along $(v_{1},v_{2})$, depending on whether $t_{2}$ is in the queue $Q$ of intersected edges888that is tested in O(1) using a per-triangle array of marks., depending on ${\tt orient\_2d(i,j,v_{0})}$ and depending on whether $t_{2}$’s first vertex is $v_{1}$ or $v_{2}$:
@@
So now the question is how to implement these three functions given that some of our vertices are given as intersections between segments and triangles. The strategy here is to keep the algorithm as near as possible to “simply implement the math”. In other words, it means we are pushing most of the difficulties towards the geometric kernel (predicates and constructions) to keep the overall structure of the algorithm simple (nearly a verbatim copy of the textbook algorithm). So we are going to compute the coordinates of the intersection points explicitely. Since these coordinates are rational fractions, and since computations for $x,y$ and $z$ are inter-related, it is reasonable to represent the intersection points in homogeneous coordinates $\hat{{\bf p}}=[x\ y\ z\ w]$ that corresponds to a 3D point ${\bf p}=[x/w\ y/w\ z/w]$. Each individual $x,y,z,w$ coordinate is represented in an exact number type that exactly implements addition, subtraction and multiplication. This is a natural representation for intersections, see (Edelsbrunner and Mücke, 1994) for an excellent tutorial on the topic, see also CGAL Homogeneous Kernel, or recent works such as (Nehring-Wirxel et al., 2021). We shall see later (§2.4 p.2.4) two alternatives to implement exact number types, and how to implement all the predicates that we need.
The input of the constrained Delaunay triangulation is a triangle $t_{1}$ and the list of segments to be inserted into $t_{1}$. Each segment’s extremity is encoded symbolically, as a triple $\sigma_{1},t_{2},\sigma_{2}$, indicating the location of the intersection within $t_{1}$, the other facet $t_{2}$ and the location of the intersection within $t_{2}$. Note that there can be also individual points (for instance, when a vertex is exactly located in a facet). The first task to do is computing the coordinates of each intersection. Depending on the nature of $\sigma_{1}$ and $\sigma_{2}$, there are three different cases, plus an additional case for intersecting constraints (four cases in total):
@@
Let ${\bf q}_{1}$ and ${\bf q}_{2}$ denote the extremities of the edge and let ${\bf p}_{1}$,${\bf p}_{2}$ and ${\bf p}_{3}$ denote the vertices of the triangle. We are in 3D if ${\bf q}_{1}$ and ${\bf q}_{2}$ are not both in the supporting plane of ${\bf p}_{1}$,${\bf p}_{2},{\bf p}_{3}$ (in other words, at least one of orient_3d$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf q}_{1})$ and orient_3d$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf q}_{2})$ is non-zero). Note that if we are in 3D, the edge $e_{1}$cap edge $e_{2}$ configuration can be seen as a particular case of the edge $e_{1}$cap triangle $t_{2}$ case, by replacing $e_{2}$ by the triangle $t_{2}$ it comes from. At the previous step, we already determined that the intersection exists, so to compute its coordinates we just need to compute a line-plane intersection, without needing to check that the intersection is in the segment and in the triangle. The intersection point ${\bf I}$ belongs to the line: